# Enabling Efficient Optimal Control of Reconfigurable Battery Systems by Unifying System Modeling and Narrowing Search Space

### v1

@@ III-B Formulation of the unified RBS model
In the above state vector, $V=[v_{1},v_{2},\dots,v_{N}]^{T}$ is the vector of battery cell OCVs, and $I_{*}=[i^{*}_{1},i^{*}_{2},\ldots,i^{*}_{N}]^{T}$ denotes the current vector of all components specified by the wildcard *. For instance, $I_{B}$, $I_{S_{m}}$, and $I_{R_{1}}$ represent the current vectors of all battery cells, all the $m$-th switches around battery cells, all $R_{1}$ in the battery cell ECMs, respectively. In addition, the system’s total current $i_{t}$ is regarded as one input. The system’s output can be selected according to the objective function $J$.
For each battery cell, the current $i_{n}^{B}={\mathrm{d}Q_{n}}/{\mathrm{d}t}$, $n=1,2,\ldots,N_{B}$, where $Q_{n}$ denotes the cell’s amount of charge. Each cell’s SoC $z_{n}={Q_{n}}/{Q_{C,n}}$, where $Q_{C,n}$ denotes its charge capacity. The slope of the cell’s OCV-SoC curve at $z_{n}$ is defined by $k_{n}^{V}(z)={\mathrm{d}v_{n}}/{\mathrm{d}z_{n}}$. Then, the evaluation of each cell’s OCV is described by
$\dot{v}_{n}=\frac{\mathrm{d}v_{n}}{\mathrm{d}t}=\frac{\mathrm{d}v_{n}}{\mathrm{d}z_{n}}\frac{\mathrm{d}z_{n}}{\mathrm{d}t}=\frac{k_{n}^{V}(z)}{Q_{C,n}}\frac{\mathrm{d}Q_{n}}{\mathrm{d}t}=\frac{k_{n}^{V}(z)}{Q_{C,n}}i_{n}^{B},$ (4)
@@
To derive the state-space equations describing the evolution of $I_{R_{1}}$ and $I_{R_{2}}$ in the state vector $X$, KCL and KVL equations need to be formulated based on the FSC in Fig. 3, where $i^{S_{m}}_{n}$ denotes the current of the $m$-th switch around the $n$-th cell, $m=1,2,\ldots,5$, $n=1,2,\ldots,N_{B}$. The KCL equations at the positive and negative electrodes of the n-th cell are formulated as follows.
@@
$\sum_{n=1}^{N}i^{S_{5}}_{n}=i_{t},$ (8)
$\sum_{n=1}^{N}i^{S_{3}}_{n}=i_{t}.$ (9)
For the four meshes between the $n$-th and $(n+1)$-th cells, $1\leq n<N_{B}$, the KVL equations are formulated as follows.
@@
By substituting the $I_{S_{5}}$ in (17) with the $I_{S_{5}}$ derived from (16), we can obtain (20). Similarly, the $I_{S_{3}}$ in (18) can be substituted with the one obtained from (19), leading to (21).
@@
Finally, based on the defined state vector $X$ in (3), we can reorganize all state evolution equations in (5), (24), and (25) to formulate the state-space model or plant for the battery system as follows.
@@
In order to pursue higher computational efficiency and match the discrete-time system plant in the formulated optimal control problem at the beginning of Section III, the state-space model in (26) can be further discretized according to, e.g., the Euler method. During each time step, battery ECM parameters might change with battery states and temperatures, and, thus, they need to be updated within the time step, leading to a time-varying system model.
Based on the system plant formulated in (26), we can derive the expression for any system output of interest, such as the cell current vector $I_{B}$, cell terminal voltage vector denoted by $V_{B}$, and the system’s terminal voltage denoted by $v_{t}$ as shown in Fig. 3. $I_{B}$ is already expressed in (22) in terms of the system’s state vector $X$ and input $i_{t}$. Denote the terminal voltage of the $n$-th battery cell by $v^{B}_{n}$, $n=1,2,\ldots,N_{B}$. Then, the cell terminal voltage vector is $V_{B}=[v^{B}_{1},v^{B}_{2},\ldots,v^{B}_{N_{B}}]^{T}$. According to the battery cell ECM in Fig. 3,
$v^{B}_{n}$ $=v_{n}-R_{0,n}i_{n}^{B}-R_{1,n}i^{R_{1}}_{n}-R_{2,n}i^{R_{2}}_{n},$ (27)
$v_{t}$ $=v^{B}_{n}-R^{S_{3}}_{n}i^{S_{3}}_{n}-R^{S_{5}}_{n}i^{S_{5}}_{n}.$ (28)
@@
Due to the instability of the terminal voltage in RBSs, the input current is influenced not only by the load power but also by the system configuration. Here, we present a simple and accurate method for determining the input current, with its accuracy demonstrated in Section III-C. Based on (30), given a power request $p_{t}$ during each time step in practical applications of battery systems, the system’s input current $i_{t}$ can be calculated by solving the quadratic equation
@@
During the above formulation of a unified RBS state-space model, it is worth noting that, the specified SSV for each time step determines the switch resistances in the KVL equations (10)-(13) through the proposed switch resistance variable in (2), and, hence, affects the system matrices in the system plant (26) and the output matrices in the output equations, such as (22), (29), and (30). This indicates that the evolution of system state variables and output variables can be controlled through adjusting the SSV. As previously mentioned, since the RBS design (a) encompasses other designs (b)–(f) shown in Fig. 1, their system plant can be constructed by selectively fixing the corresponding switch states in SSV to 0.
@@ III-C Model validation by simulation and experimental tests
To evaluate the simulation accuracy of the proposed RBS modeling framework during the operation as well as dynamic reconfiguration, a three-cell RBS was designed based on the structure in Fig. 1 (d), consisting of three cells and seven switches indexed from 1 to 7. The RBS was discharged with a constant load power of 8 W. During the discharging process, the battery system was sequentially reconfigured every 20 seconds to each configuration described by the SSV in Table I.
TABLE I: 12 feasible configurations of the RBS described by SSVs.
@@
To conduct the discharge test of the RBS, a testbed was set up, as shown in Fig. 4. Particularly, a drive circuit was designed for the MOSFET switch, and they were integrated on a printed circuit board (PCB) according to the RBS structure design in Fig. 1 (d). The RBS was discharged with a constant power by a direct current (DC) load. A data acquisition module is used to collect the voltage and current measurement data of battery cells at a sampling frequency of 10 Hz. Based on the collected measurement data, a desktop computer and a National Instruments (NI) controller are deployed to generate control signals and send them to the drive circuit of switches.
@@
In addition to the system-level simulation, the simulated cell-level currents and voltages obtained from a simple model and the proposed model are compared with experimental results in Figures 6 and 7, respectively. The simple model shares the same battery parameter values with the proposed model, but the total current is always assumed to be equally divided among parallel branches by neglecting the resistances of switches and wires in the simple model. According to the simulation MAEs for the proposed and simple models compared in Table II, we can see that the cell-level simulation results from the proposed model have current MAEs $\leq 20$ mA and voltage MAEs $\leq 7$ mV, much lower than those from the commonly used simple model assuming even parallel current distribution. The low simulation MAEs of the proposed model are mainly attributed to its capability of accurately characterizing the time-varying uneven current distribution among parallel cells during time intervals such as 80$\sim$160 seconds and 220$\sim$240 seconds in Fig. 6, corresponding to Config. 5 to Config. 8 and Config. 12 in Table I, respectively. In such configurations with parallel connections, it is of great importance to capture the dynamic current distribution for improving the state estimation accuracy and keeping the safe system operation.
@@
IV Construction of narrowed search space
After addressing the critical challenge of formulating a unified RBS plant in Section III, we are still confronted with the issue of low search efficiency when solving the formulated optimal control problems since the search space $\mathcal{B}^{N_{S}}$ grows exponentially with the increasing number of switches $N_{S}$. In this section, we will develop a framework to narrow the search space for high efficiency as well as safe connections. At first, the mapping relation between a system configuration and a system SSV will be established. Then, an algorithm will be developed to enumerate all feasible system configurations. Finally, based on the established mapping relation, only those feasible system configurations will be converted to feasible SSVs to construct a feasible search space $\mathcal{B}_{f}^{N_{S}}$. By tailoring the search scope from all system configurations to only those feasible ones, the size of the constructed feasible search space is expected to be much smaller than that of the complete search space. Since the RBS design (a) in Fig. 1 can be turned into other designs (b)-(f) by always connecting certain switches, it is taken as an example for introduce how to narrow the search space.
@@ IV-A Mapping a system configuration with a system SSV
In a battery system, all battery cells connected in series (or parallel) in a local area form an internally series-connected (or parallel-connected) battery module, referred to as a series (or parallel) module, denoted by $M_{s}$ (or $M_{p}$). If one single cell does not fall into any series or parallel module, it is regarded as a $M_{s}$ (or $M_{p}$) if cells are connected in series (or parallel) in its neighboring module. Similarly, all local battery modules can construct a series or parallel battery pack. Through analyzing the feasible configurations of the RBS design (a), we can see that some battery cells can be selected to form only series modules, and all series modules can then constitute a parallel pack. This is referred to as a series-then-parallel configuration, denoted by $C_{sp}$. On the other hand, battery cells can be selected to form some series and parallel modules, and all modules can then constitute a series pack. Such a configuration is referred to a parallel-then-series configuration, denoted by $C_{ps}$. However, at most one battery pack can be constructed among all feasible configurations of the RBS design (a).
In such a hierarchical description of a battery pack, each battery cell belongs to one unique associated module, and further one unique associated pack. Once a battery cell’s position and connection pattern (e.g., series, parallel, or single) within its associated module as well as its associate module’s position and connection pattern (P&CP) within its associated pack are specified, the state vector of all associated switches with the cell, referred to as a cell SSV, will be determined. All cell SSVs compose the RBS’s SSV, and, thus, describes the system’s configuration. For instance, for the circled cell in the RBS design (a) in Fig. 1, its cell SSV consists of the states of switches 1 to 5. The P&CPs of a battery cell and its associated module are referred to as cell-module P&CPs.
All possible cell-module P&CPs are summarized in Table III, and the corresponding cell SSVs are listed and indexed. These cell SSVs are illustrated in Fig. 8 as well. As shown in Table III, different cell-module P&CPs can be described by the same cell SSV. Moreover, if multiple cell SSVs lead to the same cell-module P&CPs, only one of them is reserved in Table III . For instance, a cell can get connected to the next cell in parallel to construct a parallel module by connecting switches 4 and 5 (i.e., cell SSV (4)), switches 1 and 4 (i.e., cell SSV (5)), or switches 1 and 3 (i.e., cell SSV (6)). Among these three cell SSVs, the first and third ones can facilitate bypassing cells but they need the top or bottom bus lines for linking the two battery system terminals. Thus, cell SSVs (4) and (6) are deployed to construct the parallel modules at the beginning and end of a configuration, respectively.
TABLE III: Cell SSVs for all possible cell-module P&CPs in the RBS design (a) in Fig. 1.
@@
Figure 8: Examples of all eight possible cell connection patterns in the RBS design (a) in Fig. 1.
Thus, given any configuration of a battery pack, we can identify each cell’s cell-module P&CPs and look up its corresponding cell SSV based on Table III. All the obtained cell SSVs are then used to construct the system SSV. Next, in order to formulate the feasible search space of system SSVs we need to identify all those feasible system configurations.
@@ IV-B Identifying all feasible system configurations
As introduced in Section I, the RBS features a wide range of terminal voltages, which can be divided into multiple levels. For an RBS with $N_{B}$ cells, the system’s terminal voltage level $L_{v_{t}}$ is set to $n_{s}\in\{1,\ldots,N_{B}\}$ if we have $n_{s}$ series-connected battery cells along the connection path between the system’s positive and negative terminals. Among all configurations of an $N_{B}$-cell RBS, there exist a majority of infeasible system configurations resulting in short circuit faults, unacceptable terminal voltage levels, or undesired open circuits. An algorithm will be developed to identify the remaining feasible configurations, only having safe connections and meeting the required terminal voltage level $L^{req}_{v_{t}}\in\{L_{v_{t}}^{min},\ldots,L_{v_{t}}^{max}\}$.
In a series-then-parallel configuration $C_{sp}$, all series-connected cells in each series module $M_{s}$ along with those bypassed cells in between are neighboring cells, and all $M_{s}$’s need to meet the required system terminal voltage level $L^{req}_{v_{t}}$, indicating that each $M_{s}$ must have $L^{req}_{v_{t}}$ cells. Moreover, any cell in a $C_{sp}$ can be bypassed. Denote all battery cells with indices from 1 to $N_{B}$, the $j$-th series module by $M_{s,j}$, and the total number of series modules by $N_{M_{s}}$. Then, a $C_{sp}$ can be constructed as follows by sequentially selecting $L^{req}_{v_{t}}$ battery cells for each series module while skipping those bypassed cells. As a result, a total of $\sum\limits_{m=2}^{\lfloor\frac{N_{B}}{L^{req}_{v_{t}}}\rfloor}\binom{mL^{req}_{v_{t}}}{N_{B}}$ feasible $C_{sp}$’s are obtained, in which $\binom{mL^{req}_{v_{t}}}{N_{B}}$ calculates a binomial coefficient.
$M_{s,j}=(k_{j,1},k_{j,2},\dots,k_{j,L^{req}_{v_{t}}}),$
$1\leq k_{j,1}<k_{j,2}<\ldots<k_{j,L^{req}_{v_{t}}}<k_{j+1,1}\leq N_{B},$
@@
$N_{M_{s}},j\in\mathbb{Z},~{}1<N_{M_{s}}\leq\frac{N_{B}}{L^{req}_{v_{t}}},~{}j\leq N_{M_{s}}.$
@@
Different from the above series-then-parallel configurations allowing only series modules with the same number of cells and arbitrary cell bypassing, a parallel-then-series configuration $C_{ps}$’s allows both series and parallel modules, and the number of cells in each module can differ from each other. Moreover, in the $C_{ps}$’s of the RBS design (a), all cells can be bypassed except those in a $M_{p}$ at the middle of a $C_{ps}$. Consequently, it becomes more difficult to identify all feasible $C_{ps}$’s. To address this challenge, Algorithm 1 is developed to identify all $C_{ps}$’s using dynamic programming. As shown in Algorithm 1, the cell-module P&CPs listed in Table III are further aggregated, and five types of position flags of battery cells are defined, including the cell bypassing. Then, a vector consisting of all cell position flags is used to describe a $C_{ps}$. To accommodate all cell flag vectors, a three-dimension list $D$ is constructed. Next, dynamic programming (DP) is deployed to search for the cell flag vectors corresponding to all feasible $C_{ps}$. DP creates sets iteratively, leveraging information from previously established sets to optimize brute-force enumeration. By storing and reusing intermediate results, it avoids redundant calculations and significantly reduces time complexity.
Algorithm 1 Algorithm of parallel-then-series Configuration Enumeration

### v2

@@ III-B Formulation of an RBS model based on its CSC
In the above state vector, $V=[v_{1},v_{2},\dots,v_{N_{B}}]^{T}$ is the vector of battery cell OCVs, and $I_{*}=[i^{*}_{1},i^{*}_{2},\ldots,i^{*}_{N_{B}}]^{T}$ denotes the current vector of all components specified by the wildcard *. For instance, $I_{B}$, $I_{S_{m}}$, and $I_{R_{1}}$ represent the current vectors of all battery cells, all the $m$-th switches around battery cells, and all $R_{1}$’s in the battery cell ECMs, respectively. In addition, the system’s total current $i_{t}$ is regarded as the system disturbance. The system’s output can be selected according to the objective function $J$.
For the $n$-th battery cell, the current $i_{n}^{B}={\mathrm{d}Q_{n}}/{\mathrm{d}t}$, $n=1,2,\ldots,N_{B}$, where $Q_{n}$ denotes the cell’s amount of charge. The $n$-th cell’s SoC is represented with $z_{n}={Q_{n}}/{Q_{C,n}}$, where $Q_{C,n}$ denotes its charge capacity. According to the $n$-th cell’s OCV-SoC curve, the $n$-th cell’s OCV denoted by $v_{n}$ can be regarded as a function of cell SoC $z_{n}$, and hence the slope of $v_{n}$ at any specific $z_{n}$ is defined by $k_{n}^{V}(z_{n})=\frac{\mathrm{d}v_{n}}{\mathrm{d}z_{n}}$. Then, the state-space equation for each cell’s OCV is described by
$\dot{v}_{n}=\frac{\mathrm{d}v_{n}}{\mathrm{d}t}=\frac{\mathrm{d}v_{n}}{\mathrm{d}z_{n}}\frac{\mathrm{d}z_{n}}{\mathrm{d}t}=\frac{k_{n}^{V}(z_{n})}{Q_{C,n}}\frac{\mathrm{d}Q_{n}}{\mathrm{d}t}=\frac{k_{n}^{V}(z_{n})}{Q_{C,n}}i_{n}^{B},$ (4)
@@
To derive the state-space equations describing the evolution of $I_{R_{1}}$ and $I_{R_{2}}$ in the state vector $X$, KCL and KVL equations need to be formulated based on the CSC in Fig. 3, where $i^{S_{m}}_{n}$ denotes the current of the $m$-th switch around the $n$-th cell, $m=1,2,\ldots,5$, $n=1,2,\ldots,N_{B}$. The KCL equations at the positive and negative electrodes of the n-th cell are formulated as follows.
@@
$\sum_{n=1}^{N_{B}}i^{S_{5}}_{n}=i_{t},$ (8)
$\sum_{n=1}^{N_{B}}i^{S_{3}}_{n}=i_{t}.$ (9)
For the four meshes between the $n$-th and $(n+1)$-th cell, $1\leq n<N_{B}$, the KVL equations are formulated as follows.
@@
By substituting the $I_{S_{5}}$ in (17) with the $I_{S_{5}}$ derived from (16), we can obtain (20). Similarly, the $I_{S_{3}}$ in (18) can be substituted with the one obtained from (19), leading to (21). In equations (20) and (21), $J_{*}$ and $K_{*}$ are computed from $E_{*},F_{*},G_{*},$ and $H_{*}$.
@@
Finally, based on the defined state vector $X$ in (3), we can reorganize all state evolution equations in (5), (24), and (25) to formulate the state-space model for the battery system as follows.
@@
In order to pursue higher computational efficiency and match the discrete-time system model in the formulated optimal control problem at the beginning of Section III, the state-space model in (26) can be further discretized according to, e.g., Zero-Order Holder method or Euler method. Considering that some ECM parameter values and the OCV-SoC slope $k_{n}^{V}$ of battery cells vary with cell SoCs, they are updated based on the new SoCs and the pre-identified look-up tables at the beginning of every sampling interval. Then, the coefficient matrices $A$, $B$, $C$, and $D$ in the system model dependent on these cell parameter values are updated accordingly. Thus, battery parameter values and coefficient matrices could change within a decision step consisting of multiple sampling intervals.
Based on the system model formulated in (26), we can derive the expression for any system output of interest, such as the cell current vector $I_{B}$, cell terminal voltage vector denoted by $V_{B}$, and the system’s terminal voltage denoted by $v_{t}$ as shown in Fig. 3. $I_{B}$ is already expressed in (22) in terms of the system’s state vector $X$ and input $i_{t}$. Denote the terminal voltage of the $n$-th battery cell by $v^{B}_{n}$, $n=1,2,\ldots,N_{B}$. Then, the cell terminal voltage vector is $V_{B}=[v^{B}_{1},v^{B}_{2},\ldots,v^{B}_{N_{B}}]^{T}$. According to the battery cell ECM in Fig. 3,
$v^{B}_{n}$ $=v_{n}-R_{0,n}i_{n}^{B}-R_{1,n}i^{R_{1}}_{n}-R_{2,n}i^{R_{2}}_{n}.$ (27)
@@
Due to the variability of the terminal voltage in RBSs, the input current is influenced not only by the load power but also by the system configuration. Here, we present a simple and accurate method for determining the input current, with its accuracy demonstrated in Section III-C. Based on (30), given a power request $p_{t}$ during each sampling interval in practical applications of battery systems, the system’s input current $i_{t}$ can be calculated by solving the quadratic equation
@@
During the above formulation of a unified RBS state-space model, it is worth noting that the specified SSV for each sampling interval determines the switch resistances in the KVL equations (10)-(13) through the proposed switch resistance variable in (2), and, hence, affects the system matrices in the system model (26) and the output matrices in the output equations, such as (22), (28), and (30). This indicates that the evolution of system state and output variables can be controlled by adjusting the SSV.
As mentioned earlier, RBS design (a) in Fig. 1 can be adapted to realize all other designs (b) to (f) along with their hybrid combinations. Therefore, the above system model formulated for design (a) is also applicable to designs (b) to (f) by selectively fixing the relevant switch states to zero in their SSVs. More importantly, the unified modeling framework introduced in this study can be deployed to model any RBS of interest, providing a solid foundation for the simulation, analysis, control, and optimization of RBSs.
@@ III-C Model validation by simulation and experimental tests
To evaluate the simulation accuracy of the proposed RBS modeling framework during operation and dynamic reconfiguration, a three-cell RBS was designed based on the structure in Fig. 1 (d), consisting of three cells and seven switches indexed from 1 to 7. The RBS was discharged with a constant load power of 8 W. During the discharging process, the battery system was sequentially reconfigured every 20 seconds to each configuration described by the SSV in Table I. Those configurations leading to short circuits or open circuits are ignored.
TABLE I: SSVs for 12 configurations of the RBS without incurring battery faults.
@@
To conduct the discharge test of the RBS, a testbed was set up, as shown in Fig. 4. Particularly, a drive circuit was designed for the MOSFET switch, and they were integrated on a printed circuit board (PCB) powered through a DC-DC converter according to the RBS structure design in Fig. 1 (d). The RBS was discharged with a constant power by a direct current (DC) load. A data acquisition module is used to collect the voltage and current measurement data of battery cells at a sampling frequency of 10 Hz. A desktop computer and a National Instruments (NI) controller are deployed to generate control signals and send them to the drive circuits of switches.
@@
In addition to the system-level simulation, the simulated cell-level currents and voltages obtained from a simple model and the proposed model are compared with experimental results in Figures 6 and 7, respectively. The simple model shares the same battery parameter values with the proposed model, but the total current is always assumed to be equally divided among parallel branches by neglecting the resistances of switches and wires in the simple model. According to the simulation MAEs for the proposed and simple models compared in Table II, we can see that the cell-level simulation results from the proposed model have current MAEs $\leq 20$ mA and voltage MAEs $\leq 7$ mV, much lower than those from the commonly used simple model assuming even parallel current distribution. The low simulation MAEs of the proposed model are mainly attributed to its capability of accurately characterizing the time-varying uneven current distribution among parallel cells during time intervals such as 80$\sim$160 seconds and 220$\sim$240 seconds in Fig. 6, corresponding to Config. 5 to Config. 8 and Config. 12 in Table I, respectively. In such configurations with parallel connections, capturing the dynamic current distribution is crucial for improving state estimation accuracy and keeping safe system operation.
@@
IV Construction of feasible solution space
After addressing the critical challenge of formulating a unified RBS model in Section III, we are still confronted with the issue of low search efficiency when solving the formulated optimal control problems since the search space $\mathcal{B}^{N_{S}}$ grows exponentially with the increasing number of switches $N_{S}$. In this section, we will develop a framework to narrow the search space for high efficiency as well as safe connections. At first, the mapping relation between a system configuration and a system SSV will be established. Then, an algorithm is developed to enumerate all feasible system configurations. Finally, based on the established mapping relation, only those feasible system configurations will be converted to feasible SSVs to construct a feasible search space $\mathcal{B}_{f}^{N_{S}}$. By tailoring the search scope from all system configurations to only those feasible ones, the size of the constructed feasible search space is expected to be much smaller than that of the complete search space. Since RBS design (a) in Fig. 1 can be turned into other designs (b)-(f) by always connecting certain switches, it is taken as an example for elaboration on how to narrow the search space.
@@ IV-A Mapping a system configuration with a system SSV
In a battery system, all battery cells connected in series (or parallel) in a local area form an internally series-connected (or parallel-connected) battery module, referred to as a series (or parallel) module, denoted by $M_{s}$ (or $M_{p}$). If one single cell does not fall into any series or parallel module, it is regarded as a $M_{s}$ (or $M_{p}$) if cells are connected in series (or parallel) in its neighboring module. Similarly, all local battery modules can construct a series or parallel battery pack. Through analyzing all feasible configurations of RBS design (a), we can see that at most one battery pack can be constructed. If battery cells are selected to form only series modules, then all series modules can constitute a parallel pack. This is referred to as a series-then-parallel configuration, denoted by $C_{sp}$. On the other hand, battery cells can be selected to form some series and parallel modules, and all modules can then be connected in series to constitute a series pack. Such a configuration is referred to a parallel-then-series configuration, denoted by $C_{ps}$.
In such a hierarchical description of a battery pack, each battery cell belongs to one unique associated module, and further one unique associated pack. Once a battery cell’s position and connection pattern (P&CP), e.g., series or parallel, within its associated module as well as its associate module’s P&CP within its associated pack are specified, the states of all associated switches with the cell, referred to as a cell SSV, will be determined. For instance, for the circled cell in RBS design (a) in Fig. 1, its cell SSV consists of the states of switches 1 to 5. All cell SSVs further compose the RBS’s SSV to describe the system’s configuration. The P&CPs of a battery cell and its associated module are referred to as cell-module P&CPs.
All possible cell-module P&CPs are summarized in Table III, and the corresponding cell SSVs are listed and indexed. These cell SSVs are illustrated in Fig. 8 as well. As shown in Table III, multiple cell-module P&CPs can be described by the same cell SSV, e.g., cell SSV (1). Moreover, if multiple cell SSVs lead to the same cell-module P&CPs, only one of them is reserved in Table III . For instance, a cell can get connected to the next cell in parallel to construct a parallel module by connecting switches 4 and 5 (i.e., cell SSV (4)), switches 1 and 4 (i.e., cell SSV (5)), or switches 1 and 3 (i.e., cell SSV (6)). Among these three cell SSVs, the first and third ones can facilitate bypassing cells but they need the top or bottom bus lines for linking the two battery system terminals. Thus, cell SSVs (4) and (6) are assigned to construct the $M_{p}$’s at the beginning and end of a $C_{ps}$, respectively, and cell SSV (5) is used for constructing a $M_{p}$ at the middle of a $C_{ps}$.
TABLE III: Cell SSVs for all possible cell-module P&CPs in RBS design (a) in Fig. 1.
@@
Figure 8: Examples of all eight possible cell connection patterns in RBS design (a) in Fig. 1.
Thus, given any configuration of the RBS design (a), we can first identify each cell’s cell-module P&CPs and then look up its corresponding cell SSV based on Table III. All the obtained cell SSVs are finally used to construct the system SSV. Next, in order to formulate the feasible search space of system SSVs we need to identify all those feasible system configurations.
@@ IV-B Identifying all feasible system configurations
Among all configurations of an $N_{B}$-cell RBS, there exist a majority of infeasible system configurations resulting in short circuit faults, unacceptable open circuits, or undesired system voltages. Thus, algorithms are developed to identify the remaining feasible configurations.
To construct a feasible system configuration meeting the required system voltage, it is necessary to specify the number of series-connected battery cells along the connection path between the system’s positive and negative terminals, which is referred to as the normalized system voltage and denoted by $\tilde{v}_{t}$. For an RBS with $N_{B}$ cells, the integer-value normalized system voltage $\tilde{v}_{t}\in\{1,\ldots,N_{B}\}$. Given the required system voltage $v_{t}^{req}\in[v_{t}^{min},v_{t}^{max}]$, the required normalized system voltage $\tilde{v}_{t}^{req}$ needs to be specified accordingly to facilitate the configuration construction, i.e., $\tilde{v}_{t}^{req}\in\{\tilde{v}_{t}^{min},\dots,\tilde{v}_{t}^{max}\}$.
The lower and upper limits of the normalized system voltage, $\tilde{v}_{t}^{min}$ and $\tilde{v}_{t}^{max}$, can be generated from the original system voltage limits, $v_{t}^{min}$ and $v_{t}^{max}$, in different ways. According to (27), each battery cell’s terminal voltage is not constant but varies with multiple factors, e.g., its OCV and current. Denote the nominal voltage and the maximum and minimum allowable terminal voltages of a battery cell by $v^{B}_{nom}$, $v^{B}_{max}$, and $v^{B}_{min}$, respectively. To cover various operating scenarios when constructing the feasible configurations, $\tilde{v}_{t}^{min}$ and $\tilde{v}_{t}^{max}$ can be roughly derived based on the battery cell’s nominal voltage, i.e., $\tilde{v}_{t}^{min}=\lfloor\frac{v_{t}^{min}}{v^{B}_{nom}}\rfloor$ and $\tilde{v}_{t}^{max}=\lceil\frac{v_{t}^{max}}{v^{B}_{nom}}\rceil$. To accommodate all potentially feasible configurations under certain operating scenarios, the selection range of $\tilde{v}_{t}$ can be further extended to, e.g., $\{\lfloor\frac{v_{t}^{min}}{v^{B}_{min}}\rfloor,\ldots,\lceil\frac{v_{t}^{max}}{v^{B}_{max}}\rceil\}$. Through these ways, some configurations might fail to meet the required system voltage under certain operating scenarios, but they can be further screened by setting relevant operating constraints when formulating control or optimization problems.
In a series-then-parallel configuration $C_{sp}$, all series-connected cells along with those bypassed cells in between in each series module $M_{s}$ are neighboring cells. If the $C_{sp}$ needs to meet the required normalized system voltage $\tilde{v}_{t}^{req}$, each $M_{s}$ in the $C_{sp}$ must have $\tilde{v}_{t}^{req}$ series-connected cells. For RBS design (a), any cell in a $C_{sp}$ can be bypassed. In a $C_{sp}$ of an $N_{B}$-cell RBS with $N_{M_{s}}$ parallel modules, $N_{M_{s}}\in[2,\lfloor\frac{N_{B}}{\tilde{v}_{t}^{req}}\rfloor]$, denote the $j$-th series module by $M_{s,j}$, $1\leq j\leq N_{M_{s}}$, the $k$-th series-connected cell in $M_{s,j}$ by $c_{j,k}$, $1\leq k\leq\tilde{v}_{t}^{req}$, and the index of cell $c_{j,k}$ among all cells in the RBS by $l_{j,k}\in[1,N_{B}]$. Then, a $C_{sp}$ can be constructed in terms of cell indices as follows by sequentially selecting $\tilde{v}_{t}^{req}$ battery cells for each $M_{s}$ while skipping those bypassed cells. As a result, a total of $\sum\limits_{N_{M_{s}}=2}^{\lfloor\frac{N_{B}}{\tilde{v}_{t}^{req}}\rfloor}\binom{N_{M_{s}}\tilde{v}_{t}^{req}}{N_{B}}$ feasible $C_{sp}$’s are obtained, in which $\binom{m\tilde{v}_{t}^{req}}{N_{B}}$ calculates a binomial coefficient.
@@
$M_{s,j}=(l_{j,1},l_{j,2},\dots,l_{j,\tilde{v}_{t}^{req}}),$
$1\leq l_{j,1}<l_{j,2}<\ldots<l_{j,\tilde{v}_{t}^{req}}<l_{j+1,1}\leq N_{B},$
$N_{M_{s}},j\in\mathbb{Z},~{}1<N_{M_{s}}\leq\frac{N_{B}}{\tilde{v}_{t}^{req}},~{}1\leq j\leq N_{M_{s}}.$
@@
Different from the above series-then-parallel configurations allowing only series modules with the same number of cells and arbitrary cell bypassing, a parallel-then-series configuration $C_{ps}$ allows both series and parallel modules, and the number of cells in each module can differ from each other. Moreover, in a $C_{ps}$ of RBS design (a), all cells can be bypassed except those in a $M_{p}$ at the middle of the $C_{ps}$. Consequently, it becomes more difficult to identify all feasible $C_{ps}$’s. To address this challenge, Algorithm 1 is developed to identify all feasible $C_{ps}$’s using dynamic programming (DP), which creates sets iteratively, leveraging information from previously established sets to optimize brute-force enumeration. By storing and reusing intermediate results to avoid redundant calculations, the search time can be significantly reduced.
Algorithm 1 Algorithm for enumerating all feasible parallel-then-series configurations